OpenStudy (anonymous):

simplify : 3/36xy x 2/6xy

OpenStudy (lncognlto):

So, just to clarify, \[\frac{ 3 }{ 36 } xy \times \frac{ 2 }{ 6 }xy\]

OpenStudy (anonymous):

yes

OpenStudy (lncognlto):

We simplify the fractions, resulting in 3/36 becoming 1/12 and 2/6 becoming 1/3, and then we multiply them together.

OpenStudy (mathmale):

Incognito is wise to clarify this. If at all possible, corenthea, would you post your problems using the Equation Editor (as Incognito has done)?

OpenStudy (lncognlto):

Therefore,\[\frac{ 1 }{ 12 } \times \frac{ 1 }{ 3 } = ?\] and \[xy \times xy = ?\]

OpenStudy (anonymous):

thank you incognlto :)

OpenStudy (lncognlto):

You're welcome. :) Hope this helped...